# When AI Evaluates Its Own Work: Validating Learner-Initiated, AI-Generated Physics Practice Problems

### v1

@@ Abstract
Large language models (LLMs) can now generate physics practice problems in real time, yet the educational value of these items hinges on rapid, reliable post-generation vetting. We investigated which automated checks are both technically feasible and pedagogically meaningful when exercises are produced on demand within a chatbot interface. A cohort of 34 introductory-physics students generated and attempted 543 problems during exam preparation. Each item was labeled by an expert on a wide range of quality attributes and presented to the learners in pairs to record their preference. We then (i) benchmarked three commodity LLMs as “judges” against the expert labels, (ii) quantified which attributes predict student choice via random-forest models, and (iii) triangulated these results with free-form exit surveys. Only a small subset of the original rubric proved necessary to reliably address student preferences either directly or by proxy. The study demonstrates that scalable formative assessment does not require exhaustive scoring: a carefully curated core of structural and learner-visible checks is sufficient to ensure both technical soundness and user appeal. The findings provide a practical blueprint for deploying real-time, AI-generated practice in physics and other quantitative disciplines.
@@ I Introduction
I.1 Background
@@
Problem solving is an essential component of physics teaching and learning [1, 2, 3]. Besides summative assessments like quizzes and exams, formative assessments such as homework and exercise sheets also play a central role in learning physics [4, 5]. To provide rapid feedback, interactive online homework has proven effective in learning physics across a number of studies [6, 7, 8, 9, 10, 11]. These materials, however, are instructor-provided and may not address the individual needs and desires of the learners. Students oftentimes ask instructors for more such practice problems before exams.
@@
Artificial intelligence (AI) is increasingly being embraced in physics education [12, 13], offering promising applications for both teaching and learning physics. Large language models (LLMs) have proven valuable for instructors, helping in the creation of tailored materials and tasks [14, 15, 13], providing personalized feedback to students [16, 17], assisting in grading [18, 19], and training of teaching staff [20]. Students report that these models represent always-available resources, which appear to be non-judgmental, imperturbable, and ever-patient [21, 22, 23],
@@
While AI-models have shown remarkable performance in solving physics problems in a number of studies [24, 25, 26, 27], generating such assessment materials is an active area of research. Despite their potential, the quality of these automatically generated exercises varies widely [28, 29], sometimes leading to student confusion and misconceptions [30]. Thus, in most current studies, instructors use the AI system to generate the questions, and then the instructors themselves judge which generated problems they present to the students [31].
@@
To enable students to generate high-quality problems on demand from course materials, we explore automated quality metrics using LLM-as-a-judge methodologies. In contrast to prior studies that relied on synthetic data, we specifically evaluate exercises produced during real interactions between students and an LLM-powered chatbot. Figures 1 and 2 show examples, which also highlight how contextual information can effectively be integrated via Retrieval-Augmented Generation (RAG) techniques [32].
@@
Figure 1: Two interactive problems generated on-demand, showcasing how course-specific contextual information, in this case the premise of the problem “with the outer wall of the apartment building,” is incorporated via Retrieval Augmented Generation (RAG) [32].
@@
I.2 Formative assessment quality
Formative assessment problems for physics are most effective when carefully designed to target key concepts and common difficulties. The quality of each question and its answer choices is crucial for measuring understanding and promoting learning [33]. As it turns out, generation of problems is one of the most frequent uses of AI by instructors [13]. In the context of independent study or small-group exam preparation, well-crafted problems (including solution paths [34]) can help students identify their own ideas about physics, practice problem-solving skills, and relate physics to the real world [35, 36, 37].
In multiple-choice questions, the incorrect options (“distractors”) should be written to be both plausible and diagnostic. Instead of obvious throw-away wrong answers, effective distractors often reflect common alternative or intuitive conceptions that students hold about the topic [38]. For example, a question on Newton’s laws might include a distractor aligning with the intuitive conception that a force is needed to keep an object moving [39]. Using known learner ideas as distractors serves two purposes: (1) it makes the question challenging and discriminating (students who have the respective alternative conception will be tempted by that option, whereas those with a firm understanding will not), and (2) it provides learners with insight into their own thinking.
@@ I.2 Formative assessment quality
Context-rich physics problems embed familiar, realistic scenarios, such as driving a car or playing sports, to tap into students’ prior experiences, intuition [40], or interests to increase motivation [42]. Unlike abstract exercises, these problems often include extraneous or irrelevant data, requiring students to discern what information is actually needed. This design mirrors real-world situations, discourages rote formula use, and encourages deeper processing: students must translate the scenario into a solvable physics problem, fostering transferable problem-solving skills. Such richly contextualized problems may have the potential to improve engagement, comprehension, and motivation by making physics feel more meaningful and relevant [40, 43]. Although more challenging, they can promote discussion and collaborative strategy-building, as they are often designed to be too difficult to solve mechanically alone but manageable in small groups, thus reflecting authentic scientific practice [40]. In this way, context-rich problems not only make learning more engaging but also develop students’ ability to identify pertinent information and apply knowledge flexibly.
@@
Each of the features above contributes to making formative assessment questions more effective learning tools. Targeted distractors ensure that a multiple-choice question probes the student’s conceptual understanding, rather than factual recall, by eliciting and addressing known misconceptions. Thoughtful numerical values and problem settings prevent students from bypassing understanding through rote computation; instead, they promote estimation, unit sense, and recognition of what a reasonable answer should be. Authentic contexts increase engagement and help students connect physics concepts to the world around them, which can improve knowledge retention and the ability to transfer learning to new situations. Formative assessment is most beneficial when it actively involves students in thinking like experts: checking their reasoning, confronting misconceptions, and making sense of results [44].
I.3 Automated exercise generation
@@
Historically, generating situational formative assessments on-the-fly and on-demand required experienced human tutors or instructors [45]. Automated exercise generation relied on rule-based algorithms, predefined templates, and ontology-driven methods, restricting flexibility [46, 47, 48]. Natural-language processing (NLP)-driven approaches, such as those designed to teach languages [49], and transformer-based systems, such as those designed to generate programming exercises [50], improved flexibility but remained limited by rigid structural dependencies.
@@
Recent advancements in large language models (LLMs) have opened new possibilities for automated generation and evaluation of exercises. However, current approaches often lack robust automated frameworks to filter out low-quality exercises [14], use synthetic datasets insufficiently capturing authentic interactions [51, 52, 53, 54], and rely on limited evaluation criteria weakly correlated with expert assessments [24, 55]. To address these gaps, our study aims to work with authentic and human-verified data: a comprehensive dataset of genuine student-chatbot interactions in an exam preparation context, capturing authentic student behavior, and a systematic evaluation of automated metrics against expert and student judgments, highlighting solution correctness, clarity, and cognitive complexity as key quality predictors. These contributions provide a foundation for improved reliability and effectiveness in AI-assisted education.
A major property of generative AI is that it always generates something. Just like in grading assessments, solutions need to be improved from both ends:
  • Improving the output of the generative AI system [56, 19]
  • Improving the evaluation of the quality of the output [18].
The objective of our study is to address the latter.
I.4 Study objective
@@
The goal of our study is to find quality metrics for the validation of exercises generated by generative AI which are reliable, relevant, and automatically assessable. The goal of this validation is to distinguish viable exercises that can be forwarded to the learners from defective exercises that should be discarded.
@@
We are using a dataset of exercises generated by students in an authentic study setting, and we establish ground truth on a set of potentially useful and usable quality metrics through an expert assessment; in other words, we consider the expert assessment to be the “truth” on which we base the remainder of the study. We then establish the reliability of automatically assessing these metrics using an LLM-as-a-judge, and we establish relevance by correlating these quality metrics with measures of user preferences.
@@ II.1 Target system
Figure 3: Information flow in an enhanced chatbot that can generate verified problems on-the-fly and on-demand. The “exercise validation” step (green box) is at the subject of our study.
@@
Once validated, the JSON data representing selected exercises is transmitted to the user’s browser, where JavaScript dynamically renders interactive problems based on this structured data. Client-side scripts evaluate student-submitted answers for correctness (see Fig. 2). Students can continue their dialogue with the chatbot or revisit and retry previously generated exercises by scrolling through the chat history.
During the dataset generation for this study, only 11-out-of-NN multiple-choice (“radio-button”) and numerical exercise types were implemented. Exercise validation at this stage involved only basic sanity checks, resulting in a dataset with varying and frequently insufficient quality.
@@ II.2 Dataset
A total of $543$ exercises were generated by the participants, where one student prompt could result in more than one exercise (see Fig. 1). Students would select the exercise candidate that they wanted to work with and attempt to solve. For performance reasons, in later laboratory sessions, the answer for numerical problems would not be generated by the server until after the student selected the candidate. For each generated exercise, the following data were collected:
@@
  • JSON code of the exercise
@@
Additional data collected included a post-survey on their experiences, which allowed for free-form entry.
II.3 Quality metrics
Table 1 shows quality metrics that are determined from the question and the preceding dialogue. Table 2 shows metrics that are determined based solely on the generated problem for both problem types; Table 3 shows additional metrics that only pertain to numerical problems, while Table 4 shows the same for 11-out-of-NN multiple-choice problems.
Table 1: Metrics derived from preceding prompts as part of the ground truth.

### v2

@@ Abstract
Large language models (LLMs) can now generate physics practice problems in real time, yet the educational value of these items hinges on rapid, reliable post-generation vetting. In this exploratory study, we investigated which automated checks are both technically feasible and pedagogically meaningful when exercises are produced on demand within a chatbot interface. A cohort of 34 introductory-physics students generated and attempted 543 practice problems during exam preparation. Each item was labeled by an expert on a wide range of quality attributes and presented to the learners in pairs to record their preference. We then (i) benchmarked three commodity LLMs as “judges” against the expert labels, (ii) quantified which attributes predict student choice via random-forest models, and (iii) triangulated these results with free-form exit surveys. Only a small subset of the original metric items proved necessary to reliably address student preferences either directly or by proxy. The study demonstrates that scalable formative assessment does not require exhaustive scoring: a carefully curated core of structural and learner-visible checks is sufficient to ensure both technical soundness and user appeal. The findings provide a practical blueprint for deploying real-time, AI-generated practice in physics and other quantitative disciplines.
@@ I Introduction
I.1 Motivation
Physics instructors routinely observe that students ask for additional practice problems on particular topics, especially in the run-up to exams. Students may have worked through end-of-chapter exercises, old exams, and other available formative assessments, yet still struggle with specific concepts and feel unprepared for upcoming summative assessments. In large-enrollment courses, instructors typically cannot respond to individualized, topic-specific requests — and they certainly cannot write new practice problems on demand in a scalable way.
In such situations, students who can afford it may turn to private tutoring. In principle, generative AI offers a more equitable alternative: students can ask a chatbot to generate practice problems on demand. In practice, however, when these systems lack access to course materials, they may produce problems that are misaligned with the course, underspecified or internally inconsistent, or accompanied by incorrect solutions. Figure 1 illustrates this failure mode using GPT-5.2 Thinking [1]. The dialogue begins with a standard, solvable problem: the scenario and questions are clearly formulated, all required information is present (and the value of $g$ is only specified in part (c), where it is first needed), and the chosen height $h$ indeed keeps the car on track. When asked for solutions, the system also computes parts (a)–(c) correctly. We would therefore consider parts (a)–(c) “valid” in a physics sense. However, the final optional part contains a serious physical error: the claimed “approximation” $N\approx mg+mv^{2}/R$ is not justifiable in this context. When asked to justify it, the system argued that the approximation is needed to make the integral easier — which is true, but we would argue that the problem is physically “invalid” with that misleading additional information.
Figure 1: An example of an on-demand problem generated by GPT-5.2 Thinking.
In short, for a variety of reasons, AI-generated problems may be unsuitable for practice even before pedagogical considerations enter the picture. An unsolvable, contradictory, or physically incorrect problem has little formative value (unless diagnosing its flaws is itself an intended learning objective [2, 3]). In this study, we therefore investigate metrics that can be used to verify the “nuts-and-bolts” validity of on-demand generated physics practice problems. Using problems generated by student prompt during a simulated study session as samples, we evaluate candidate validation metrics by their agreement with expert judgment and by their relationship to surface-level student expectations.
I.2 Background
@@
Problem solving is an essential component of physics teaching and learning [4, 5, 6]. Besides summative assessments like quizzes and exams, formative assessments such as homework and exercise sheets also play a central role in learning physics [7, 8]. To provide rapid feedback, interactive online homework has proven effective in learning physics across a number of studies [9, 10, 11, 12, 13, 14]. These materials, however, are instructor-provided and may not address the individual needs and desires of the learners (an alternative are student-authored problems, which may vary in quality [15]).
@@
Artificial intelligence (AI) is increasingly being embraced in physics education [16, 17, 18], offering promising applications for both teaching and learning physics. Due to growing corpora of training materials and improved reasoning algorithms, Large language models (LLMs) have developed strong conceptual reasoning capabilities [19] (even though their conceptual framework might not always be stable [20]). They have thus, proven valuable for instructors, helping in the creation of tailored materials and tasks [21, 22, 17], providing personalized feedback to students [23, 24], assisting in grading [25, 26], and training of teaching staff [27]. Students report that these models represent always-available resources, which appear to be non-judgmental, imperturbable, and ever-patient [28, 29, 30].
@@
While AI-models have shown remarkable performance in solving physics problems in a number of studies [31, 32, 33, 19], generating such assessment materials is an active area of research. Despite their potential, the quality of these automatically generated problems varies widely [34, 35], sometimes leading to student confusion and misconceptions [36]. Thus, in most current studies, instructors use the AI system to generate the problems, and then the instructors themselves judge which generated problems they present to the students [37].
@@
To enable students to generate valid problems on demand from course materials, we explore automated metrics using LLM-as-a-judge methodologies [38]. In contrast to prior studies that relied on synthetic data, we specifically evaluate problems produced during real interactions between students and an LLM-powered chatbot. Figures 2 and 3 show examples, which also highlight how contextual information can effectively be integrated via Retrieval-Augmented Generation (RAG) techniques [39].
@@
Figure 2: Two interactive problems generated on-demand, showcasing how course-specific contextual information, in this case the premise of the problem “with the outer wall of the apartment building,” is incorporated via Retrieval Augmented Generation (RAG) [39]. As opposed to traditional chatbot output as in Fig. 1, the problems are rendered with an interactive answer field.
@@
I.3 Formative assessment quality
Formative assessment problems for physics are most effective when carefully designed to target key concepts and common difficulties. In this subsection, we use quality in the pedagogical sense (supporting valid measurement and learning); later we separately operationalize relevance as alignment with learner uptake in the deployment setting. The quality of each problem, namely the question and its answer choices, is crucial for measuring understanding and promoting learning [40]. As it turns out, generation of problems is one of the most frequent uses of AI by instructors [17]. In the context of independent study or small-group exam preparation, well-crafted problems (including solution paths [41]) can help students identify their own ideas about physics, practice problem-solving skills, and relate physics to the real world [42, 43, 44].
In multiple-choice problems, the incorrect options (“distractors”) should be written to be both plausible and diagnostic. Instead of obvious throw-away wrong answers, effective distractors often reflect common alternative or intuitive conceptions that students hold about the topic [45]. For example, a problem on Newton’s laws might include a distractor aligning with the intuitive conception that a force is needed to keep an object moving [46]. Using known learner ideas as distractors serves two purposes: (1) it makes the problem challenging and discriminating (students who have the respective alternative conception will be tempted by that option, whereas those with a firm understanding will not), and (2) it provides learners with insight into their own thinking.
@@ I.3 Formative assessment quality
Context-rich physics problems embed familiar, realistic scenarios, such as driving a car or playing sports, to tap into students’ prior experiences, intuition [47], or interests to increase motivation [49]. Unlike abstract problems, these problems often include extraneous or irrelevant data, requiring students to discern what information is actually needed. This design mirrors real-world situations, discourages rote formula use, and encourages deeper processing: students must translate the scenario into a solvable physics problem, fostering transferable problem-solving skills. Such richly contextualized problems may have the potential to improve engagement, comprehension, and motivation by making physics feel more meaningful and relevant [47, 50]. Although more challenging, they can promote discussion and collaborative strategy-building, as they are often designed to be too difficult to solve mechanically alone but manageable in small groups, thus reflecting authentic scientific practice [47]. In this way, context-rich problems not only make learning more engaging but also develop students’ ability to identify pertinent information and apply knowledge flexibly.
@@
Each of the features above contributes to making formative assessment problems more effective learning tools. Targeted distractors ensure that a multiple-choice problem probes the student’s conceptual understanding, rather than factual recall, by eliciting and addressing known misconceptions. Thoughtful numerical values and problem settings prevent students from bypassing understanding through rote computation; instead, they promote estimation, unit sense, and recognition of what a reasonable answer should be. Authentic contexts increase engagement and help students connect physics concepts to the world around them, which can improve knowledge retention and the ability to transfer learning to new situations. Formative assessment is most beneficial when it actively involves students in thinking like experts: checking their reasoning, confronting misconceptions, and making sense of results [51].
I.4 Automated problem generation
@@
Historically, generating situational formative assessments on-the-fly and on-demand required experienced human tutors or instructors [52]. Automated problem generation relied on rule-based algorithms, predefined templates, and ontology-driven methods, restricting flexibility [53, 54, 55]. Natural-language processing (NLP)-driven approaches, such as those designed to teach languages [56], and transformer-based systems, such as those designed to generate programming exercises [57], improved flexibility but remained limited by rigid structural dependencies.
@@
Recent advancements in large language models (LLMs) have opened new possibilities for automated generation and evaluation of problems. However, current approaches often lack robust automated frameworks to filter out low-quality problems [21], use synthetic datasets insufficiently capturing authentic interactions [58, 59, 60, 61], and rely on limited evaluation criteria weakly correlated with expert assessments [31, 62]. To address these gaps, our study aims to work with authentic and human-verified data: a comprehensive dataset of genuine student-chatbot interactions in an exam preparation context, capturing authentic student behavior, and a systematic evaluation of automated metrics against expert and student judgments, highlighting solution correctness, clarity, and cognitive complexity as key quality predictors. These contributions provide a foundation for improved reliability and effectiveness in AI-assisted education.
A core characteristic of generative AI is that it always produces some output—regardless of its relevance, accuracy, or grounding. As a result, LLM-generated problems may be unclear, misleading, or entirely fabricated (“hallucinations”). Ensuring reliable, high quality automated problem generation, therefore, requires addressing both ends of the process:
  1. 1.
    Improving the output of generative AI systems [63, 26]
  2. 2.
    Improving the evaluation of the quality of the output [25]
The former is akin to an instructor drafting problems for an assignment or an exam and represents a broad research area; the latter parallels the “proofreading” done by colleagues or teaching assistants. In this study, we focus exclusively on the latter.
I.5 Study objective
@@
The goal of our study is to identify validation metrics for problems generated by generative AI that are reliable, relevant, and automatically assessable. We employ an LLM-as-a-judge mechanism, where the initial LLM output is assessed by the LLM using a set of metrics. The goal of this validation is to distinguish viable problems that can be forwarded to the learners from defective problems that should be discarded.
@@
We are using a dataset of problems generated on-demand by our LLM-powered chatbot during student-chatbot interactions in a user study emulating exam preparation. The system generated a problem whenever it detected that a student was requesting one; see Figure 2 for an example prompt. We establish ground truth on a set of potentially useful and usable quality metrics through an expert rating process; in other words, we consider these human expert ratings to be the “truth” on which we base the remainder of the study. We then establish the reliability of automatically assessing these same metrics using an LLM-as-a-judge, and we establish relevance by correlating these quality metrics with measures of user preferences.
I.6 Operational definitions and research questions
We distinguish between pedagogical quality (an item’s potential to diagnose understanding and support learning; see Sect. I.3), technical defects, and relevance to learner uptake. While ideal validation would relate item properties to independent efficacy measures such as IRT-based item parameters or observed learning gains, those require repeated administrations or longitudinal designs that are outside the scope of this study and not available in our on-demand, conversational practice context. We therefore operationalize quality-related constructs within the constraints of our data and deployment setting.
Within the framework of our study, the terms reliable and automatically assessable refer to properties of the evaluation procedure, while relevance refers to observable student choice (not necessarily learning efficacy):
Reliable
Defined as model–ground truth agreement. A metric is called reliable if an LLM-as-a-judge can reproduce the expert label with high agreement.
Relevant
Defined as predictive of learner choice and preference. A metric is called relevant if it is associated with what students actually prefer when selecting between candidate items. Our system use case is that students are shown multiple candidate AI-generated problems and must choose what to attempt; therefore, a necessary condition for any pedagogically strong item to matter in practice is that it is actually selected and engaged with. We accordingly interpret relevance as perceived usefulness/appeal and practical fit (triangulated with exit-survey feedback), rather than as a direct measure of learning efficacy.
Automatically assessable
Defined as feasible at scale without human intervention. A metric is automatically assessable if it can be computed end-to-end by an LLM from the generated problem (JSON) and its immediate chat context, returning a schema-conformant JSON judgment with no human intervention, external solution key, or additional retrieval beyond the conversation.
These definitions lead to the following research questions:
RQ1 (Reliability)
Which quality metrics can commodity LLMs judge with strong agreement to expert labels on confirmatory samples, and how does this depend on model choice?
RQ2 (Relevance)
Which metrics are predictive of student choice between candidate problems for (a) multiple-choice and (b) numerical items, and how do these findings triangulate with free-form exit-survey feedback?
RQ3 (Automatic assessability)
Within the intersection of reliable and relevant metrics, which subset can be automatically assessed by an LLM fast and cheaply enough for real-time deployment, and how small can this subset be without materially degrading predictive power?
@@ II Methodology
The following subsections describe the implementation of our study. Note that in several of the technical implementation details, for example the labels of the metrics and the prompts, practice problems were denoted as exercises and in rare cases as questions — we have reproduced these in the original.
@@ II.1 Target system
Figure 4: Information flow in an enhanced chatbot that can generate verified practice problems on-the-fly and on-demand. The “problem validation” step (LLM-as-a-judge; green box) is the subject of our study.
@@
Once validated, the JSON data representing selected problems is transmitted to the user’s browser, where JavaScript dynamically renders interactive problems based on this structured data. Client-side scripts evaluate student-submitted answers for correctness (see Fig. 3). Students can continue their dialogue with the chatbot or revisit and retry previously generated problems by scrolling through the chat history.
During the dataset generation for this study, only 11-out-of-NN multiple-choice (“radio-button”) and numerical problem types were implemented. Problem validation at this stage involved only basic sanity checks, resulting in a dataset with varying and frequently insufficient quality.
@@ II.2 Dataset
Typical examples of student prompts which triggered problem generation were, translated from German, are:
  • Please give me an entry-level problem about torque.
  • Give me an exam problem about mechanical energy.
  • Please give me a problem involving Kirchhoff rules.
  • Give me a conceptual question about kinematics.
  • Give me complex problem about electric fields.
  • Give me a more difficult problem about kinematics.
  • Another one.
For prompts like the last one, “another one,” the intent detection concluded that the user desired another problem like the most recently generated one. Generally, problem prompts were not very specific when it came to desired scenarios or solution methods. Examples of general prompts, again translated from German, are:
  • How do I find the sign of the torque?
  • Please explain the solution path for this problem.
  • What are 0.08 mJ in J?
  • Why is this formula only valid for a straight path?
A total of $543$ problems were generated in response to the participants, where one student prompt could result in more than one problem (see Fig. 2). Students would select the problem candidate that they wanted to work with and attempt to solve. For performance reasons, in later laboratory sessions, the answer for numerical problems would not be generated by the server until after the student selected the candidate. For each generated problem, the following data were collected:
@@
  • JSON code of the problem
@@
Additional data collected included a post-survey on their experiences, which allowed for free-form entry. It asks about perceived usefulness and whether the generated problems met expectations, as well as the quality, clarity, and trustworthiness of the chatbot’s answers and explanations. It also probes perceived learning efficiency compared to traditional methods, effects on motivation, and how well the chatbot adapted to individual learning needs. The Likert-type questions (translated from German to English) were:
  • How useful was the chatbot for your preparation for the mini-exam?
  • How useful did you find the automatic generation of practice problems?
  • Did the automatic generation of practice problems meet your expectations?
  • How do you rate the quality of the answers generated by the chatbot?
  • How clear were the chatbot’s explanations?
  • How trustworthy did you find the chatbot’s information?
  • How efficient was learning with the chatbot compared to traditional methods?
  • How did using the chatbot affect your motivation to learn?
  • How well was the chatbot able to address your individual learning needs?
The survey also included an open-ended field for comments, remarks, etc.
II.3 A note about nomenclature
In this manuscript, we use problem to mean a typical exam or homework item consisting of a question prompt and an associated answer (e.g., a solution or answer choices). In the analysis code and data structures created earlier in the project (repository listed under Data Availability), the terms problem , task, and exercise appear as legacy variable names and are used interchangeably; for transparency and reproducibility, we retain those identifiers in the shared code and tables provided here, but we use the standardized terminology in the remainder of the manuscript. Finally, when students used the German term Fragen, we translate it as questions, even when the context indicates they were referring to complete problems as defined above.
II.4 Validity metrics
The metrics we used for validating the generated problems were inspired by work on metrics for problem solutions [65]; we asked ourselves what properties a problem would need to have to demand solutions that address those solution metrics. We added metrics that assess the implicit quality criteria of solvability, completeness, and correctness of the expected answer. Both sets of metrics align with earlier work on assessing the quality of student-generated problems [15]. To our knowledge, the resulting set is the largest set assembled thus far in literature, however, we still cannot claim completeness; we thus designed the validation mechanism for possible expansion in future studies.
Table 1 shows quality metrics that are determined from the problem and the preceding dialogue. Table 2 shows metrics that are determined based solely on the generated problem for both problem types; Table 3 shows additional metrics that only pertain to numerical problems, while Table 4 shows the same for 11-out-of-NN multiple-choice problems. Here, “N/A” is an explicit ground-truth label indicating that the desired level was not specified or inferable from the chat context (rather than an unscored item). For the agreement analyses in Sect. III.1 we treat N/A as an additional class for these intent metrics; their low macro-level scores therefore reflect that the intended Bloom/difficulty level is often not recoverable from the available context. For the relevance models in Sect. III.2, we excluded these two predictors (rather than excluding the corresponding items), so the remaining analyses are unaffected except that we do not make claims about the influence of these intent-level constructs on student choice.
Table 1: Metrics derived from preceding prompts as part of the ground truth. In our technical implementation, the terms exercise and (practice) problem were used interchangeably.